It took MegaTech, Inc. 200,000 labor-hours to produce the first of several oil drilling rigs for Antarctic exploration. Your company, Natural Resources, Inc. has agreed to purchase the fifth oil drilling rig from their manufacturing yard. Assume that MegaTech experiences a learning rate of 75%. At a labor rate of $38 per hour, what should you, as the purchasing agent expect to pay for the fifth unit? TN = T1C, Where TN = Time needed to produce the Nth unit

Rate, labor/hr

$38

learning rate

0.75

T1 = Time needed to produce the first unit

200,000

C5th =learning curve coefficient given, look for it in table.

TN

Total cost

a.

TN is $102,600

b.

Total cost is $3,898,800

c.

C5th =learning curve coefficient given, look for it in table is 0.513

alternative without table. It took MegaTech, Inc. 100,000 labor-hours to produce the first of several oil drilling rigs for Antarctic exploration. Your company, Natural Resources, Inc. has agreed to purchase the fifth oil drilling rig from their manufacturing yard. Assume that MegaTech experiences a learning rate of 80%. At a labor rate of $35 per hour, How long should all five oil-drilling rigs take to complete? TN = T1C, Where TN = Time needed to produce the Nth unit

labor hour

$35

learning rate

0.8

a = the time required for the initial unit of output

100,000

X, the Nth unit

1

2

3

4

5

a. b = the slope of the learning curve, represented as: log decimal learning rate/log 2

b. TNx, Time needed to produce the Nth unit

c. TN cumulative

d. total costs

a.

b = the slope of the learning curve, represented as: log decimal learning rate/log 2 is -0.3219

b.

TNx, Time needed to produce the third unit is less than 64,000

c.

TN cumulative time for all units is $373,774

d.

total costs is $13,082,094

*Multiple Answers May Be Selected For Each Question*
Answer #1

1st question

learning curve coefficient given C = 0.513 for learning rate 75% and for 5th unit

So, Tn = 200000*0.513 = 102600

Total cost is 102600*38 = 3898800

all the statements a. b. and c. are correct


Alternative without table

Formula

Correct answers are a.

b = the slope of the learning curve, represented as: log decimal learning rate/log 2 is -0.3219 = LOG(0.8)/LOG(2) = -0.3219

c.TN cumulative time for all units is $373,774

d.total costs is $13,082,094

b. is incorrect as time taken to produce 3rd unit is 70120 hours